### v3

@@ 1 Introduction
Common graphs that are not bipartite, i.e., their chromatic number is larger than two, are scarce. In particular, Jagger, Šťovíček and Thomason asked whether there exists a common graph with chromatic number at least four. While odd cycles and even wheels [27, 40] are examples of $3$-chromatic common graphs, also see [22], the existence of a common graph with chromatic number at least four was open until 2012 when the $5$-wheel was shown to be common [25] as one of the first application of the flag algebra method of Razborov [35]. The question whether there exist common graphs with arbitrarily large chromatic number has been reiterated in [25], and also by Conlon, Fox and Sudakov in the survey paper “Recent developments in graph Ramsey theory”, which is now a classical reference for results and directions related to Ramsey theory.
Problem (Conlon, Fox and Sudakov [8, Problem 2.28]).
@@
We treat common graphs using methods from the theory of graph limits, which allows us applying spectral tools from the operator theory. To prove Theorem 1, we show that every graph of sufficiently large girth can be embedded in a graph $H$ such that any $2$-edge-coloring of a complete graph has asymptotically at least as many monochromatic copies of $H$ as the random $2$-edge-coloring. The proof of Theorem 1 is split into two cases based on whether the considered $2$-edge-coloring is close to the random coloring or not; we refer to the two cases as the local regime and the non-local regime (we provide an overview of the proof in Subsection 1.2). The core of the proof is formed by the arguments related to the local regime, which is described by the existence of a dominant eigenvalue of the operator associated with the $2$-edge-coloring. While both the Sidorenko property and common graphs have been studied in the local regime [32, 19, 18, 13, 23], the proof of Theorem 1 requires new spectral arguments to control (in)dependence of monochromatic embeddings of different parts of $H$ in the host edge-colored complete graph. Our techniques could be used in other settings, e.g., the setting of a conjecture of Kohayakawa, Nagle, Rödl and Schacht [29] as discussed in Section 7.
@@ 1.1 Proof motivation
However, if a graphon $W$ contains a sparse part, it may even happen that $t(H,W)=0$, so $t(H\cup K_{n,n},W)=0$, and the argument fails. In fact, it can be shown that the presence of a sparse part is essentially the only obstacle for the argument to proceed. However, if $W$ contains a sparse part of measure $\alpha_{0}>0$ with density $q$ close to 00, then this part has density $1-q$ in $1-W$ and its contribution to $t(H\cup K_{n,n},1-W)$ (as formalized in Lemma 21) is roughly $\alpha_{0}^{v\left(H\right)+2n}(1-q)^{e\left(H\right)+n^{2}}$; the value of $\alpha_{0}$ depends on $H$ only and so if $n$ is sufficiently large, then
@@ 1.2 Proof overview
$W$ with density $p\geq p_{0}$ and $\|W-p\|_{\square}\leq\varepsilon$ normal-$\Rightarrow$ $t(H_{0},W)\geq p^{e(H_{0})}$
@@
$W$ with density $p\geq p_{0}$, no sparse part and $\|W-p\|_{\square}\geq\varepsilon$ normal-$\Rightarrow$ $t(H_{0},W)\geq p^{e(H_{0})}$
@@
As the first step towards proving Theorems 23 and 24, we deal with the local regime and establish a counterpart of Theorem 3, which is given as Theorem 18 in Section 3. While Theorem 18 applies to graphs of a particular form only, the bound on the cut distance given in Theorem 3 is independent of the length $\ell$ of the path attached to the graph $H$ (in addition to relaxing the technical assumption that $\|W\|_{\infty}\leq 2p$ as discussed in Subsection 1.1). The independence on $\ell$ is essential for proving Theorem 19 in Section 4, where the obtained bound on the cut distance have to be independent of the parameters $\ell$, $m$ and $n$ so that it is sufficiently strong to eventually yield proofs of Theorems 23 and 24. The proof of Theorem 19, which is the technically most challenging part of our argument, is obtained using spectral arguments, which allow us to control how densities of rooted subgraphs change when walking along a path in a graphon. In particular, we argue that the density of $H_{0}$ in a graphon $W$ is at least the density of $H_{0}$ in the constant graphon with the same density as $W$ (Cases (i) and (ii) in the proof) unless the graphon $W$ has a sparse part (Case (iii)), which in turn implies that it also has a denser part where sufficiently many copies of the graph $H_{0}$ can be found.
@@
Finally, Section 6 presents our main results—Theorem 23 and 24; the main steps of their proofs are visualized in Figure 1. If a graphon $W$ or $1-W$ has a sparse part, then the density of $H_{0}$ in the complement of the sparse part is sufficiently large to establish that $H_{0}$ is common; here, we need that $\ell$ is sublinear in $v(H)+mn$. Otherwise, if a graphon $W$ is close to a constant graphon, then the density of $H_{0}$ in both $W$ and $1-W$ is at least the density of $H_{0}$ in the constant graphons with the same density by Theorem 19, and if $W$ is not close to a constant graphon (and does not have a sparse part), then the density of $H_{0}$ in both $W$ and $1-W$ is at least the density of $H_{0}$ in the constant graphons with the same density by Theorem 22. A simple convexity argument now yields that $t(H_{0},W)+t(H_{0},1-W)\geq 2\cdot 2^{-e(H_{0})}$.
The proof of the more general Theorem 24 proceeds by induction on $k$, which is the number of colors. Individual color classes are represented by graphons $W_{1},\ldots,W_{k}$ such that $W_{1}+\cdots+W_{k}$ is a graphon with density close to one (rather than exactly equal to one); this makes the induction argument significantly easier. If one of graphons $W_{1},\ldots,W_{k}$ has a very sparse part, we apply induction to the remaining $k-1$ colors restricted to the sparse part. If one of the color classes is sparse, we argue that we find sufficiently many copies of $H_{0}$ in the densest color class (here, we use that $\ell$ is sublinear in $v(H)+mn$). If neither of these two cases apply, we use for each graphon $W_{i}$, $i=1,\ldots,k$, Theorem 19 or Theorem 22, depending whether $W_{i}$ is close to a constant graphon or not, i.e., whether the corresponding color class falls into the local or non-local regime, to derive that the density of $H_{0}$ in each $W_{i}$ is at least as in the constant graphon with the same density, and we eventually derive using convexity of the function $p^{e(H_{0})}$ that the sum of densities of $H_{0}$ in $W_{1},\ldots,W_{k}$ is sufficiently large.
@@ 2.1 Graph limits
where the supremum is taken over all measurable subsets $S$ and $T$ of $[0,1]$. If $U$ is a kernel with $\|U\|_{\infty}\leq 1$, then it holds [33, Lemma 8.12] that
@@
Lemma 4 (Cooper, Kráľ and Martins [12, Lemma 11]).
@@
On the other hand, the $H$-Counting Lemma, which we now state, implies that $t(C_{4},W)\leq p^{4}+4\|W-p\|_{\square}$ for every graphon $W$ with density $p$.
Lemma 5 ($H$-Counting Lemma, Lovász [33, Lemma 10.23]).
@@ 2.3 Spectral properties of graphons and kernels
We now review spectral properties of graphons and kernels; we refer to [33, Section 7.5] for further details. Fix a kernel $U$ with $\|U\|_{\infty}\leq 1$. We can think of $U$ as a Hilbert-Schmidt integral operator from $L_{2}[0,1]$ to $L_{2}[0,1]$ defined as
@@
$U(x,y)=\sum_{i}\lambda_{i}f_{i}(x)f_{i}(y)$
for almost every $(x,y)\in[0,1]^{2}$. We remark that if $U$ is a graphon with density $p$, then $\lambda_{1}$ is at least $p$. Since we have assumed that $\|U\|_{\infty}\leq 1$, it holds that $(Uf_{i})(x)\leq\|f_{i}\|_{2}=1$ for every $x\in[0,1]$, i.e., $\|Uf_{i}\|_{\infty}\leq 1$, which implies that $\|f_{i}\|_{\infty}\leq\lvert\lambda_{i}\rvert^{-1}$.
@@
To express the density of paths, we need to introduce additional notation. Let ${\mathfrak{j}}:[0,1]\to[0,1]$ be the constant function equal to $1$, and set $\alpha_{i}\in[0,\pi/2]$ to be the real such that $\langle{\mathfrak{j}},f_{i}\rangle=\cos\alpha_{i}$ (we may assume without loss of generality by replacing $f_{i}$ with $-f_{i}$ that the product is non-negative). Further set $\delta=1-\cos\alpha_{1}$. Informally speaking, for a graphon $W$, $\delta$ measures how close $W$ is to having all the degrees the same, in particular, $\delta=0$ if and only if $\deg_{W}(x)=p$ for almost every $x\in[0,1]$ where $p$ is the density of $W$. The following holds for every $n\geq 2$:
@@
In the rest of this subsection, we deal with graphons only and estimate some of the introduced parameters. These estimates will be used repeatedly in the proofs of Theorems 18 and 19. Fix a graphon $W$ with density $p\in[0,1]$, and let $\gamma=t(C_{4},W)-p^{4}$. Note that $\gamma\geq 0$ since the cycle $C_{4}$ has the Sidorenko property; on the other hand, $\gamma\leq 4\|W-p\|_{\square}$ by the $H$-Counting Lemma (Lemma 5). Since the density of $C_{4}$ in $W$ is at least $\lambda_{1}^{4}$, it follows that
@@ 2.4 Estimates on densities
We now review several estimates on densities of graphs in graphons and kernels. A large number of the estimates that we present is standard and we state them explicitly just for referencing in the rest of the paper. We start with one of such estimates, which follows by the standard use of Jensen’s inequality, see e.g. [30, Proposition 1.10]. The two lemmas are also implied by the weak Hölder property that all complete bipartite graphs have since they are weakly norming [24], also see [33, Section 14.1] for in-depth discussion.
@@
[…]
@@
[…]
Lemma 11 (Lovász [32, Corollary 3.12]).
@@ 3 Locally Sidorenko component
In this section, we establish that a certain family of graphs that contain any fixed high girth graph as an induced subgraph has the Sidorenko property in a local sense. As discussed in Subsections 1.1 and 1.2, the main contribution of this section is Theorem 18 asserting that the value of $\varepsilon_{0}$ is independent of $\ell$ and so of the graph $H^{\bullet}\oplus P_{\ell}^{\bullet}$. The proof is inspired by the proof of an extension of a result by Lovász [32] due to Fox and the last author [19, 18], however, it does not follow from either of these results even when the bound is not required to be uniform. We remark that the lower bound of $50$ on the girth of $G$ given in Theorem 18 can certainly be decreased, possibly all the way to $4$, at the expense of making the proof of Theorem 18 significantly more technical; we decided to prove the theorem with the lower bound of $50$ as proving it with an even larger lower bound on the girth would not simplify the proof substantially.
@@
for every $m\geq 1$. The density of $H_{\ell}$ can be expressed (see [32] and [33, Proof of Proposition 16.27]) as
@@
    The graph $\langle F_{0}\rangle$ has no core component containing $v_{8}$, all core components are acyclic or $C_{4}$, and at least one of the acyclic components is not a star.
@@
We next analyze the contribution of the terms for $F_{0}$ from each of the eight groups in Claims 18.118.7. Claim 18.1 and Claim 18.7 will readily yield that the left hand side of (23) is at least
@@ 4 Local regime
Our next step is prove the following lower bound on the measure of the set $S$:
@@
Finally, we estimate $\gamma_{T}=t(C_{4},W[T])-p_{T}^{4}$ as follows:
@@
Since the density $p_{T}$ of $W[T]$ is larger than $q$ by (67) (recall that $q\leq p(1+p_{0}^{mn}/4)$) and $\gamma_{T}\leq\gamma_{0}$ by (68), the choice of $q$ implies that
@@
Theorem 19 together with the $H$-Counting Lemma (Lemma 5) applied for $H=C_{4}$ yields the following.

### v5

@@ 1 Introduction
Common graphs that are not bipartite, i.e., their chromatic number is larger than two, are scarce. In particular, Jagger, Šťovíček and Thomason asked whether there exists a common graph with chromatic number at least four. While odd cycles [42] and even wheels [28, 40] are examples of $3$-chromatic common graphs, also see [23], the existence of a common graph with chromatic number at least four was open until 2012 when the $5$-wheel was shown to be common [26] as one of the first application of the flag algebra method of Razborov [36]. The question whether there exist common graphs with arbitrarily large chromatic number has been reiterated in [26], and also by Conlon, Fox and Sudakov in the survey paper “Recent developments in graph Ramsey theory”, which is now a classical reference for results and directions related to Ramsey theory.
Problem (Conlon, Fox and Sudakov [9, Problem 2.28]).
@@
We treat common graphs using methods from the theory of graph limits while employing spectral tools from the operator theory; graph limit methods permit avoiding lower order terms when analyzing large graphs, which makes our arguments simpler to present and in some cases also more natural. To prove Theorem 1, we show that every graph of sufficiently large girth can be embedded in a graph $H$ such that any $2$-edge-coloring of a complete graph has asymptotically at least as many monochromatic copies of $H$ as the random $2$-edge-coloring. The proof of Theorem 1 is split into two cases based on whether the considered $2$-edge-coloring is close to the random coloring or not; we refer to the two cases as the local regime and the non-local regime (we provide an overview of the proof in Subsection 1.2). The core of the proof is formed by the arguments related to the local regime, which is described by the existence of a dominant eigenvalue of the operator associated with the $2$-edge-coloring. While both the Sidorenko property and common graphs have been studied in the local regime [33, 20, 19, 14, 24], the proof of Theorem 1 requires new spectral arguments to control (in)dependence of monochromatic embeddings of different parts of $H$ in the host edge-colored complete graph. Our techniques could be used in other settings, e.g., the setting of a conjecture of Kohayakawa, Nagle, Rödl and Schacht [30] as discussed in Section 7.
@@ 1.1 Proof motivation
However, if a graphon $W$ contains a sparse part, it may even happen that $t(H,W)=0$, so $t(H\cup K_{n,n},W)=0$, and the argument fails (in fact, it can be shown that the presence of a sparse part is essentially the only obstacle for the argument to proceed). However, if $W$ contains a sparse part of measure $\alpha_{0}>0$ with density $q$ close to 00, then this part has density $1-q$ in $1-W$ and its contribution to $t(H\cup K_{n,n},1-W)$ (as formalized in Lemma 21) is roughly $\alpha_{0}^{v\left(H\right)+2n}(1-q)^{e\left(H\right)+n^{2}}$; the value of $\alpha_{0}$ depends on $H$ only and so if $n$ is sufficiently large, then
@@ 1.2 Proof overview
$W$ with density $p\geq p_{0}$ and $\|W-p\|_{\square}\leq\varepsilon$Rightarrow $t(H_{0},W)\geq p^{e(H_{0})}$
@@
$W$ with density $p\geq p_{0}$, no sparse part and $\|W-p\|_{\square}\geq\varepsilon$Rightarrow $t(H_{0},W)\geq p^{e(H_{0})}$
@@
As the first step towards proving Theorems 23 and 24, we deal with the local regime and establish a counterpart of Theorem 3, which is given as Theorem 18 in Section 3. While Theorem 18 applies to graphs of a particular form only, the bound on the cut distance given in Theorem 3 is independent of the length $\ell$ of the path attached to the graph $H$ (in addition to relaxing the technical assumption that $\|W\|_{\infty}\leq 2p$ as discussed in Subsection 1.1). The independence on $\ell$ is essential for proving Theorem 19 in Section 4, where the obtained bound on the cut distance has to be independent of the parameters $\ell$, $m$ and $n$ so that it is sufficiently strong to eventually yield proofs of Theorems 23 and 24. The proof of Theorem 19, which is the technically most challenging part of our argument, is obtained using spectral arguments, which allow us to control how densities of rooted subgraphs change when walking along a path in a graphon. In particular, we argue that the density of $H_{0}$ in a graphon $W$ is at least the density of $H_{0}$ in the constant graphon with the same density as $W$ (Cases (i) and (ii) in the proof) unless the graphon $W$ has a sparse part (Case (iii)), which in turn implies that it also has a denser part where sufficiently many copies of the graph $H_{0}$ can be found.
@@
Finally, Section 6 presents our main results—Theorems 23 and 24; the main steps of their proofs are visualized in Figure 1. If a graphon $W$ or $1-W$ has a sparse part, then the density of $H_{0}$ in the complement of the sparse part is sufficiently large to establish that $H_{0}$ is common; here, we need that $\ell$ is sublinear in $v(H)+mn$. Otherwise, if a graphon $W$ is close to a constant graphon, then the density of $H_{0}$ in both $W$ and $1-W$ is at least the density of $H_{0}$ in the constant graphons with the same density by Theorem 19, and if $W$ is not close to a constant graphon (and does not have a sparse part), then the density of $H_{0}$ in both $W$ and $1-W$ is at least the density of $H_{0}$ in the constant graphons with the same density by Theorem 22. A simple convexity argument now yields that $t(H_{0},W)+t(H_{0},1-W)\geq 2\cdot 2^{-e(H_{0})}$.
The proof of the more general Theorem 24, which concern $k$-common graphs, proceeds by induction on $k$, which is the number of colors. Individual color classes are represented by graphons $W_{1},\ldots,W_{k}$ such that $W_{1}+\cdots+W_{k}$ is a graphon with density close to one (rather than exactly equal to one); this makes the induction argument significantly easier. If one of graphons $W_{1},\ldots,W_{k}$ has a very sparse part, we apply induction to the remaining $k-1$ colors restricted to the sparse part. If one of the color classes is sparse, we argue that we find sufficiently many copies of $H_{0}$ in the densest color class (here, we use that $\ell$ is sublinear in $v(H)+mn$). If neither of these two cases apply, we use for each graphon $W_{i}$, $i=1,\ldots,k$, Theorem 19 or Theorem 22, depending whether $W_{i}$ is close to a constant graphon or not, i.e., whether the corresponding color class falls into the local or non-local regime, to derive that the density of $H_{0}$ in each $W_{i}$ is at least as in the constant graphon with the same density, and we eventually derive using convexity of the function $p^{e(H_{0})}$ that the sum of densities of $H_{0}$ in $W_{1},\ldots,W_{k}$ is sufficiently large.
@@ 2.1 Graph limits
where the supremum is taken over all measurable subsets $S$ and $T$ of $[0,1]$. If $U$ is a kernel with $\|U\|_{\infty}\leq 1$, then it holds [34, Lemma 8.12] that
@@
Lemma 4 (Cooper, Kráľ and Martins [13, Lemma 11]).
@@
On the other hand, the Counting Lemma, which we now state, implies that $t(C_{4},W)\leq p^{4}+4\|W-p\|_{\square}$ for every graphon $W$ with density $p$.
Lemma 5 (Counting Lemma, Lovász [34, Lemma 10.23]).
@@ 2.3 Spectral properties of graphons and kernels
We now review spectral properties of graphons and kernels; we refer to [34, Section 7.5] for further details. Fix a kernel $U$ with $\|U\|_{\infty}\leq 1$. We can think of $U$ as a Hilbert-Schmidt integral operator from $L_{2}[0,1]$ to $L_{2}[0,1]$ defined as
@@
$\sum_{i}\lambda_{i}f_{i}(x)f_{i}(y)$
converges to $U$ in the $L_{2}$-norm. We remark that if $U$ is a graphon with density $p$, then $\lambda_{1}$ is at least $p$. Since we have assumed that $\|U\|_{\infty}\leq 1$, it holds that $(Uf_{i})(x)\leq\|f_{i}\|_{2}=1$ for every $x\in[0,1]$, i.e., $\|Uf_{i}\|_{\infty}\leq 1$, which implies that $\|f_{i}\|_{\infty}\leq\lvert\lambda_{i}\rvert^{-1}$.
@@
To express the density of paths, we need to introduce additional notation. Let ${\mathfrak{j}}:[0,1]\to[0,1]$ be the constant function equal to $1$, and set $\alpha_{i}\in[0,\pi/2]$ to be the real such that $\langle{\mathfrak{j}},f_{i}\rangle=\cos\alpha_{i}$ (we may assume without loss of generality by replacing $f_{i}$ with $-f_{i}$ that the product is non-negative). Further set $\delta=1-\cos\alpha_{1}$. Informally speaking, for a graphon $W$, $\delta$ measures how close $W$ is to having all the degrees the same, in particular, $\delta=0$ if and only if $\deg_{W}(x)=p$ for almost every $x\in[0,1]$ where $p$ is the density of $W$. The following holds for every $n\geq 2$ (for $n=2$, the equality is a particular case of (7.21) in [34]):
@@
In the rest of this subsection, we deal with graphons only and estimate some of the introduced parameters. These estimates will be used repeatedly in the proofs of Theorems 18 and 19. Fix a graphon $W$ with density $p\in[0,1]$, and let $\gamma=t(C_{4},W)-p^{4}$. Note that $\gamma\geq 0$ since the cycle $C_{4}$ has the Sidorenko property; on the other hand, $\gamma\leq 4\|W-p\|_{\square}$ by the Counting Lemma (Lemma 5). Since the density of $C_{4}$ in $W$ is at least $\lambda_{1}^{4}$, it follows that
@@ 2.4 Estimates on densities
We now review several estimates on densities of graphs in graphons and kernels. A large number of the estimates that we present is standard and we state them explicitly just for referencing in the rest of the paper. We start with one of such estimates, which follows by the standard use of Jensen’s inequality, see e.g. [31, Proposition 1.10]. The two lemmas are also implied by the weak Hölder property that all complete bipartite graphs have since they are weakly norming [25], also see [34, Section 14.1] for in-depth discussion.
@@
$\lvert t(K_{1,k},U)\rvert\leq t(K_{1,2},U)^{k/2}\leq t(K_{1,2},U)\quad\mbox{and}\quad\lvert t(K_{2,k},U)\rvert\leq t(C_{4},U)^{k/2}\leq t(C_{4},U).$
@@
$\int_{[0,1]^{k}}\prod_{i\in[n]}f_{i}(x_{1},\ldots,x_{n})\,\mbox{d}x_{1}\cdots\,\mbox{d}x_{k}\leq\prod_{i\in[n]}\|f_{i}\|_{2}.$
Lemma 11 (Lovász [33, Corollary 3.12]).
@@ 3 Locally Sidorenko component
In this section, we establish that a certain family of graphs that contain any fixed high girth graph as an induced subgraph has the Sidorenko property in a local sense. As discussed in Subsections 1.1 and 1.2, the main contribution of this section is Theorem 18 asserting that the value of $\varepsilon_{0}$ is independent of $\ell$ and so, when $H^{\bullet}$ is fixed, of the graph $H^{\bullet}\oplus P_{\ell}^{\bullet}$. The proof is inspired by the proof of an extension of a result by Lovász [33] due to Fox and the last author [20, 19], however, it does not follow from either of these results even when the bound is not required to be uniform. We remark that the lower bound of $50$ on the girth of $G$ given in Theorem 18 can certainly be decreased, possibly all the way to $4$, at the expense of making the proof of Theorem 18 significantly more technical; we decided to prove the theorem with the lower bound of $50$ as proving it with an even larger lower bound on the girth would not simplify the proof substantially.
@@
for every $m\geq 1$. The density of $H_{\ell}$ can be expressed (see [33] and [34, Proof of Proposition 16.27]) as
@@
    The graph $\langle F_{0}\rangle$ has no core component containing $v_{8}$, all core components are acyclic or $C_{4}$, and at least one of the acyclic core components is not a star.
@@
We briefly verify that each subset $F_{0}$ belongs to exactly one of the groups (a)(h). Since any core component containing the vertex $v_{8}$ can be neither $C_{4}$ nor a star, the group (d) is disjoint from the remaining seven groups. If $F_{0}$ is not covered by the group (a) or (d), then $\langle F_{0}\rangle$ has at least one core component and no core component of $\langle F_{0}\rangle$ contains $v_{8}$. We restrict our attention to such subsets $F_{0}$ in the rest of this paragraph. If $\langle F_{0}\rangle$ has a core component containing a cycle of $G$ (note that the length of this cycle is at least $50$), then $F_{0}$ belongs to the group (e) and it cannot belong to any other group. If $\langle F_{0}\rangle$ has no core component containing a cycle of $G$ but it has a core component that contains the cycle $C_{4}$ and this core component is not solely this cycle, then $F_{0}$ belongs to the group (f) and it cannot belong to any other group. Hence, we can further restrict our attention to only those subsets $F_{0}$ such that no core component of $\langle F_{0}\rangle$ contains $v_{8}$ and each core component of $\langle F_{0}\rangle$ is either acyclic or solely the cycle $C_{4}$. If one of core components is not a star, then $F_{0}$ belongs to the group (g) and it cannot belong to any other group. Otherwise, all core components of $\langle F_{0}\rangle$ are stars or solely the cycle $C_{4}$. If $F_{0}$ has a single core component, then it belongs to the group (b) or the group (c); otherwise, i.e., when $F_{0}$ has at least two core components, it belongs the group (h). We conclude that each subset $F_{0}$ belongs to exactly one of the groups (a)(h).
We next analyze the contribution of the terms for $F_{0}$ based on the groups (a)(h) in Claims 18.118.7. Claim 18.1 and Claim 18.7 will readily yield that the left hand side of (23) is at least
@@ 4 Local regime
Our next step is proving the following lower bound on the measure of the set $S$:
@@
Note that if $q$ were equal to $1$, then (67) would imply that $p_{T}>1$ (recall that $q\leq p(1+p_{0}^{mn}/4)$), which is impossible, and so it holds that $q<1$. Finally, we estimate $\gamma_{T}=t(C_{4},W[T])-p_{T}^{4}$ as follows:
@@
Since the density $p_{T}$ of $W[T]$ is larger than $q$ by (67) (again recall that $q\leq p(1+p_{0}^{mn}/4)$) and $\gamma_{T}\leq\gamma_{0}$ by (68), the choice of $q$ implies that
@@
Theorem 19 together with the Counting Lemma (Lemma 5) applied for $C_{4}$ yields the following.